# PTQ1.61: Push the Real Limit of Extremely Low-Bit Post-Training Quantization Methods for Large Language Models

### v1

@@ 3.1 Preliminaries
[…]
@@
Binarization represents the most extreme form of quantization (1-bit) where weights are assigned as ±$\pm 1$ determined by the sign function. In more details:
@@ 3.3 Block-wise Scaling Factors Optimization
Dataset Methods Bits 1-7 1-13 1-30 1-65 2-7 2-13 2-70 3-8 3-70
WikiText2 FP 16 5.68 5.09 4.10 3.53 5.47 4.88 3.31 6.14 2.93
AWQ 2 2.5e5 2.8e5 2.4e5 7.4e4 2.2e5 1.2e5 - 8.2e5 -
GPTQ 2 2.1e3 5.5e3 1.9e3 55.91 7.7e3 2.1e3 77.95 5.7e4 -
QuIP 2 42.19 12.18 9.36 7.19 55.00 13.75 6.96 119.23 51.73
OmniQuant 2 15.47 13.21 8.81 7.58 37.37 17.21 7.81 796.8 2.8e4
PB-LLM 1.7(+1) 102.19 48.11 26.37 12.91 66.30 462.84 NAN 78.67 37.99
BiLLM 1(+1.1) 35.04 15.14 10.52 8.51 32.48 21.77 12.80 50.59 89.91
PTQ1.61 1.61 12.50 9.67 7.95 7.02 12.70 9.74 6.94 22.90 28.26
C4 FP 16 7.08 6.61 5.98 5.62 6.97 6.46 5.52 8.88 6.92
AWQ 2 1.9e5 2.3e5 2.4e5 7.5e4 1.7e5 9.4e4 - 8.1e5 -
GPTQ 2 689.13 2.5e3 234.95 40.58 NAN 323.12 48.82 1.0e5 -
OmniQuant 2 24.89 18.31 13.67 10.77 90.64 26.76 12.28 2.4e3 5.5e4
PB-LLM 1.7(+1) 67.92 34.20 22.45 13.70 66.23 333.54 NAN 78.98 82.16
BiLLM 1(+1.1) 33.64 14.75 10.97 9.92 33.72 23.14 13.84 48.26 210.12
PTQ1.61 1.61 17.13 13.51 10.98 9.86 17.73 13.64 12.63 33.82 37.34
@@ 3.4 Quantization Preprocessing
It is crucial to emphasize that our preprocessing scheme stands out significant advantages and differences from existing post-quantization PEFT methods, e.g., LoRA, QLoRA and QA-LoRA (Xu et al., 2023), please refer to Appendix D.
@@ 4.1 Experimental Setup
Table 2: Reasoning accuracies comparison on LLaMA family. More tasks are listed in Table 10 and 11.
@@
We initialize learnable scaling factors with $\alpha_{w}=\frac{\lVert w\rVert_{1}}{n_{w}}$ and AdamW optimizer (Loshchilov and Hutter, 2017) with zero weight decay is utilized to update them with learning rate 5e-4 and 1e-3. For PTQ, our calibration set sampled from WikiText2 (Merity et al., 2016) consists of 128 random 2048 token-segments and the block-wise training process includes 20 epochs with a batch size of 1. For quantization preprocessing, the number of steps and ranks in lightweight restorative LoRA is 10K and 32 respectively. The entire process is deployed on 2 Nvidia A800 GPUs.
@@
On language generation tasks which are the core objectives of LLMs, our test set comes from WikiText2 and C4 (Raffel et al., 2020). We also select several reasoning benchmarks, i.e., PIQA (Bisk et al., 2020), ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), Race (Lai et al., 2017) and LAMBADA (Paperno et al., 2016), using the open-sourced toolkit lm-evaluation-harness (Gao et al., 2023). In addition, RedPajama (Computer, 2023) is used for quantization preprocessing.
@@ 4.3 Experiments on Reasoning Benchmarks
Reasoning capability is becoming a crucial metric for evaluating PTQ approaches. The comparison results are indicated in Table 2, where our method exhibits superiority in most benchmarks. For instance, compared with the second best baseline, BiLLM, our method showcases an average performance increase of $1.58\%\sim 5.92\%$. Particularly, on LAMBADA most baselines suffer from significant performance degradation while we still maintain an outstanding level. We also assess the evaluation on MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021) and LongBench (Bai et al., 2024), please refer to Appendix E. Considering the comparison results and the minimal 1.61-bit, we conclude that our method represents the state-of-the-art extremely low-bit PTQ scheme for LLMs.
@@ B.2 Distinctions with AWQ and OWQ
For AWQ, both their method and our PTQ1.61 take into account the relationship between input activation and weight, and utilize this relationship for subsequent processing. But notably, there is no structured mask in AWQ. They leverage this discovery to perform a grid search based on MSE loss to select appropriate quantization scalars. These scalars are used to scale the corresponding weights up or down. Weights corresponding to larger input activations are assigned larger scalars to increase their magnitude to reduce quantization error. Conversely, our structured mask is designed through rigorous mathematical derivation that input activation has a greater impact on the upper bound of quantization error. This allows us to assess the importance or salience of corresponding weights and avoid unacceptable binarization errors. Furthermore, the results in Table 5 demonstrate that our method outperforms AWQ in extremely low-bit quantization scenarios without requiring additional innovations.
@@ C.2 Angular Biases
$\mbox{W}_{q}^{,}=(\alpha_{r_{1}}\times\alpha_{r_{2}})\circ(\alpha_{s}\text{sign}(\mbox{W})).$ (9)
@@ Appendix D More Details on Quantization Preprocessing
D.2 Resources Requirement
@@
D.3 OPT Results on Other Approaches
@@
D.4 Comparison with Post-Quantization PEFT Methods

### v2

@@ 3.1 Preliminaries
where $\mbox{W}\in\mathbb{R}^{n\times m}$ and $\mbox{W}_{q}\in\mathbb{R}^{n\times m}$ indicate full-precision and quantized weights respectively. $\lfloor\cdot\rceil$ denotes round-to-nearest operator. $S_{q}$ is the quantization scalar and $Z_{q}$ represents zero-point.
@@
Binarization represents the most extreme form of quantization (1-bit) where weights are assigned as $\pm 1$ determined by the sign function. In more details:
@@ 3.3 Block-wise Scaling Factors Optimization
Dataset Methods Bits 1-7 1-13 1-30 1-65 2-7 2-13 2-70 3-8
WikiText2 FP 16 5.68 5.09 4.10 3.53 5.47 4.88 3.31 6.14
AWQ 2 2.5e5 2.8e5 2.4e5 7.4e4 2.2e5 1.2e5 - 8.2e5
GPTQ 2 2.1e3 5.5e3 1.9e3 55.91 7.7e3 2.1e3 77.95 5.7e4
QuIP 2 42.19 12.18 9.36 7.19 55.00 13.75 6.96 119.23
OmniQuant 2 15.47 13.21 8.81 7.58 37.37 17.21 7.81 796.8 4
PB-LLM 1.7(+1) 102.19 48.11 26.37 12.91 66.30 462.84 NAN 78.67
BiLLM 1(+1.1) 35.04 15.14 10.52 8.51 32.48 21.77 12.80 50.59
PTQ1.61 1.61 12.50 9.67 7.95 7.02 12.70 9.74 6.94 22.90
C4 FP 16 7.08 6.61 5.98 5.62 6.97 6.46 5.52 8.88
AWQ 2 1.9e5 2.3e5 2.4e5 7.5e4 1.7e5 9.4e4 - 8.1e5
GPTQ 2 689.13 2.5e3 234.95 40.58 NAN 323.12 48.82 1.0e5
OmniQuant 2 24.89 18.31 13.67 10.77 90.64 26.76 12.28 2.4e3
PB-LLM 1.7(+1) 67.92 34.20 22.45 13.70 66.23 333.54 NAN 78.98
BiLLM 1(+1.1) 33.64 14.75 10.97 9.92 33.72 23.14 13.84 48.26
PTQ1.61 1.61 17.13 13.51 10.98 9.86 17.73 13.64 12.63 33.82
@@ 3.4 Quantization Preprocessing
We emphasize that the goal of our preprocessing scheme is to produce a LLM that is more amenable to channel-wise quantization, rather than using the preprocessed model for inference directly. In Appendix D.2, we further illustrate this point with example experiments. Additionally, it is crucial to clarify that our preprocessing scheme stands out significant advantages and differences from existing post-quantization PEFT methods, e.g., LoRA, QLoRA and QA-LoRA (Xu et al., 2023), please refer to Appendix D.
@@ 4.1 Experimental Setup
Table 2: Reasoning accuracies comparison on LLaMA family. More tasks are listed in Table 11 and 12.
@@
We initialize learnable scaling factors with $\alpha_{w}=\frac{\lVert w\rVert_{1}}{n_{w}}$ and AdamW optimizer (Loshchilov and Hutter, 2017) with zero weight decay is utilized to update them with learning rate 5e-4 and 1e-3. For PTQ, our calibration set sampled from WikiText2 (Merity et al., 2016) consists of 128 random 2048 token-segments and the block-wise training process includes 20 epochs with a batch size of 1. For quantization preprocessing, the number of steps and ranks in lightweight restorative LoRA is 20K and 64 respectively. The entire process is deployed on 2 Nvidia A800 GPUs.
@@
On language generation tasks which are the core objectives of LLMs, our test set comes from WikiText2 and C4 (Raffel et al., 2020). We also select several reasoning benchmarks, i.e., PIQA (Bisk et al., 2020), ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), Winogrande (Sakaguchi et al., 2021), Race (Lai et al., 2017) and LAMBADA (Paperno et al., 2016), using the open-sourced toolkit lm-evaluation-harness (Gao et al., 2023). We also assess the evaluation on MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021) and LongBench (Bai et al., 2024), please refer to Appendix E. In addition, RedPajama (Computer, 2023) is used for quantization preprocessing.
@@ 4.3 Experiments on Reasoning Benchmarks
Reasoning capability is becoming a crucial metric for evaluating PTQ approaches. The comparison results are indicated in Table 2, where our method exhibits superiority in most benchmarks. For instance, compared with the second best baseline, BiLLM, our method showcases an average performance increase of $1.58\%\sim 5.92\%$. Particularly, on LAMBADA most baselines suffer from significant performance degradation while we still maintain an outstanding level. Considering the comparison results and the minimal 1.61-bit, we conclude that our method represents the state-of-the-art extremely low-bit PTQ scheme for LLMs.
@@ 4.5 Quantization Preprocessing on Baselines
At present, despite its strong performance, our preprocessing scheme still suffers from drawbacks such as longer runtime (as indicated in Table 8). Therefore, we must clarify that our preprocessing scheme is viewed as an optional component of our PTQ1.61. As indicated in Table 6, even without preprocessing, our PTQ1.61 still achieves state-of-the-art performance under extremely low-bit setting.
@@ 5 Conclusion
Acknowledgment
Miao Zhang was partially sponsored by the National Natural Science Foundation of China under Grant 62306084 and U23B2051, and Shenzhen College Stability Support Plan under Grant GXWD20231128102243003 and Grant ZDSYS20230626091203008.
@@ B.2 Distinctions with AWQ and OWQ
For AWQ, both their method and our PTQ1.61 take into account the relationship between input activation and weight, and utilize this relationship for subsequent processing. But notably, there is no structured mask in AWQ. They leverage this discovery to perform a grid search based on MSE loss to select appropriate quantization scalars. These scalars are used to scale the corresponding weights up or down. Weights corresponding to larger input activations are assigned larger scalars to increase their magnitude to reduce quantization error. Conversely, our structured mask is designed through rigorous mathematical derivation that input activation has a greater impact on the upper bound of quantization error. This allows us to assess the importance of corresponding weights and avoid unacceptable binarization errors. Furthermore, the results in Table 5 demonstrate that our method outperforms AWQ in extremely low-bit quantization scenarios without requiring additional innovations.
@@ C.2 Angular Biases
$\mbox{W}_{q}^{{}^{\prime}}=(\alpha_{r_{1}}\times\alpha_{r_{2}})\circ(\alpha_{s}*\text{sign}(\mbox{W})).$ (9)
@@ Appendix D More Details on Quantization Preprocessing
D.2 Preprocessed Model and FP Model
In Section 3.4, we introduce the rationale behind our quantization preprocessing scheme and emphasize its purpose is to reshape the weight distribution to be more suitable for quantization while not using the preprocessed model directly for inference. As shown in Table 9, it can be observed that after preprocessing the FP16 performance of LLaMA-13B degrades slightly, but after quantization it significantly outperforms the pretrained model, which further proves the effectiveness of our preprocessing method.
Table 9: Full-precision perplexities comparison of the pretrained and preprocessed LLaMA-13B.
Model Bits Dataset
WikiText2 C4
Pretrained 16 5.09 6.61
-Quant 1.61 14.22 20.78
Preprocessed 16 9.32 12.33
-Quant 1.61 9.67 13.51
D.3 Resources Requirement
@@
D.4 OPT Results on Other Approaches
@@
D.5 Comparison with Post-Quantization PEFT Methods